19. Subba Rao started work in 1995 at an annual salary of Rs 5000 and received an increment of Rs 200 each year. In which year did his income reach Rs 7000?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It can be seen from the given question, that the incomes of Subba Rao increases every year by Rs.200 and hence, forms an AP. Therefore, after 1995, the salaries of each year are; 5000, 5200, 5400, … Here, first term, a = 5000 and common difference, d = 200 Let after nth year, his salary be Rs 7000. Therefore, by the nth term formula of AP, an = a+(n−1) d 7000 = 5000+(n−1)200 200(n−1)= 2000 (n−1) = 10 n = 11 Therefore, in 11th year, his salary will be Rs 7000.